OpenStudy (anonymous):

Find the product. -7(-a^2)^2(-b^3) i need help, step by step.

OpenStudy (anonymous):

After you @lgbasallote

OpenStudy (lgbasallote):

accoding to PEMDAS start with exponent so what is \[(-a^2)^2?\]

OpenStudy (anonymous):

-a^4

OpenStudy (lgbasallote):

not exactly...maybe this will help \[(-2)^2 = 2^2 = 4\] because \[(-2)^2 \implies -2 \times -2\] and negative times negative is positive. does that make sense?

OpenStudy (anonymous):

oh ok, yes

OpenStudy (lgbasallote):

so what's \[(-a^2)^2\]

OpenStudy (anonymous):

four

OpenStudy (lgbasallote):

how did you get 4?

OpenStudy (anonymous):

do i do four times -7 now?

OpenStudy (lgbasallote):

no...you're doing \[(-a^2)^2\] im asking how that became 4

OpenStudy (anonymous):

oh.. a^4

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so you have \[-7(a^4)(-b^3)\] what is 7 times a^4?

OpenStudy (anonymous):

7a^4?

OpenStudy (lgbasallote):

uhh correction.. what's -7 times a^4

OpenStudy (lgbasallote):

forgot to put my negative sign

OpenStudy (anonymous):

-7a^4.

OpenStudy (lgbasallote):

right so now you have \[(-7a^4)(-b^3)\]

OpenStudy (lgbasallote):

what would be the product?

OpenStudy (anonymous):

so isit 7a^4b^3?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thank you!

OpenStudy (lgbasallote):

you're welcome!